OpenStudy (anonymous):

f(x)=-x^4-4 if i had to plug in a -2 for the x what would i get? i got 12. is that right?

OpenStudy (turingtest):

\[\large f(x)=-x^4-4\]is the problem?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

watch thh signs\[\large f(-2)=-(-2)^4-4=-(16)-4=?\]

OpenStudy (anonymous):

-20?

OpenStudy (turingtest):

looks that way :D

OpenStudy (turingtest):

did you know that (-2)^4=16 ? or did you think it was negative?

OpenStudy (anonymous):

i know it equals 16 but the negative sign infront of the x threw me off.

OpenStudy (turingtest):

right-o !

OpenStudy (anonymous):

what would it be if it was a -1 instead of a -2?

OpenStudy (turingtest):

\[(-1)^4=(-1)(-1)(-1)(-1)=(1)(1)=1\]

OpenStudy (anonymous):

ok i got that wrong too. i wrote -3

OpenStudy (anonymous):

wait, so \[-1^{4}-4\]=1?

OpenStudy (turingtest):

no, I just did the (-1)^4 part, try the rest yourself!

OpenStudy (anonymous):

-5

OpenStudy (turingtest):

yep, nice job :)

OpenStudy (anonymous):

thank you!